# Tensor Product Attention Is All You Need

### v3

@@ G.2.2 Ablation Study on Different Ranks
G.3 Ablation Studies on Learning Rates
We implement a set of parallel experiments for medium models with learning rate $3\times 10^{-4}$, and the curves for training loss, validation loss, and validation perplexity are displayed in Figure 6. We also show the performance of these models on the benchmarks described in Section 4 in Tables 12-13. The results show that TPA and TPA-KVonly models can also outperform other types of attention with different learning rates.
@@ G.3 Ablation Studies on Learning Rates
Figure 6: The training loss, validation loss, and validation perplexity of medium-size (353M) models (learning rate $3\times 10^{-4}$) and different attention mechanisms on the FineWeb-Edu 100B dataset.

### v4

@@ Appendix I More on TPA
J.3 Ablation Studies on Learning Rates
To assess sensitivity to learning rates, we conducted parallel experiments on medium-sized models using a learning rate of $3\times 10^{-4}$ (compared to the default $6\times 10^{-4}$ used for other medium model results). The training loss, validation loss, and validation perplexity curves are shown in Figure 11. Performance on standard benchmarks for these models trained with the $3\times 10^{-4}$ learning rate are reported in Tables 18 (0-shot) and 19 (2-shot). The results demonstrate that TPA and TPA-KVonly maintain their performance advantages over other attention mechanisms even with this alternative learning rate.
@@
Figure 11: The training loss, validation loss, and validation perplexity of medium-size (353M) models (learning rate $3\times 10^{-4}$) with different attention mechanisms on the FineWeb-Edu 100B dataset.